# Predicting topological invariants and unconventional superconducting pairing from density of states and machine learning

### v1

@@ I Introduction
In parallel to these developments, magnetic impurities in superconductors have been shown to generate odd-frequency pairing Kuzmanovski et al. (2020); Perrin et al. (2020); Santos et al. (2020); Suzuki et al. (2022). The study of odd-frequency superconductivity started when Berezinskii Berezinskii (1974) pointed out that the symmetry of the superconducting state allows for a two-fermion pairing function which is odd under time reversal but even under the product of spatial inversion and spin and orbital permutations. Since then, states exhibiting this type of unconventional pairing have been identified in many different systems; for a review, see Ref. Linder and Balatsky (2019). For a single magnetic impurity or dilute magnetic impurities in superconductors, the local density of states (LDOS) measured by STS can be used to estimate the imaginary part of the odd-frequency local pairing function Perrin et al. (2020); Santos et al. (2020). Remarkably, odd-frequency pairing has also been linked to the presence of Majorana bound states Daino et al. (2012); Asano and Tanaka (2013); Kashuba et al. (2017); Tsintzis et al. (2019). In particular, the anomalous Green’s function near vortices in chiral $p$-wave superconductors shows the same spatial structure as the LDOS associated with subgap Majorana modes Daino et al. (2012).
@@
In this work, we study the relation between topological invariants, edge states and odd-frequency pairing in a model for a 2D superconductor with randomly distributed magnetic impurities. Besides being naturally present in superconducting monolayers Ménard et al. (2019), the disorder can be introduced by substitution of magnetic adatoms and lattice reconstruction during the sample growth Goedecke et al. (2022) or by direct atomic manipulation Küster et al. (2022). Sufficiently strong disorder can drive transitions between topological and trivial phases Qin et al. (2016); Mascot et al. (2019). In Ref. Mascot et al. (2019), these transitions were analyzed using the Bott index. This topological invariant can be calculated in real space and is equivalent to the Chern number in the homogeneous case Hastings and Loring (2010); Loring and Hastings (2011); Hastings and Loring (2011); Toniolo (2022). The Bott index has been used to characterize topological phases in several disordered and amorphous systems Bandres et al. (2016); Agarwala and Shenoy (2017); Huang and Liu (2018); Focassio et al. (2021); Ghadimi et al. (2021); Dantas and Andrade (2022).
@@
We are interested in how the topological classification based on the Bott index correlates with the LDOS measured in STS experiments. To address this question, we use machine learning (ML) Carrasquilla (2020), a provably efficient method to classify phases of matter Huang et al. (2022); Canabarro et al. (2019). In fact, ML techniques have been successfully applied to identify topological phase transitions van Nieuwenburg et al. (2017); Deng et al. (2017); Zhang et al. (2018); Rem et al. (2019); Rodriguez-Nieva and Scheurer (2019); Scheurer and Slager (2020); Lustig et al. (2020); Claussen et al. (2020); Holanda and Griffith (2020); Che et al. (2020), to estimate model parameters from numerical or experimental data Wang et al. (2017); Gentile et al. (2021); Koch et al. (2023), and to analyze images generated by scanning probe techniques in quantum materials Burzawa et al. (2019); Iwasawa et al. (2022); Basak et al. (2023); Sobral et al. (2023). Here we train ML models to predict the Bott index from the LDOS calculated numerically for our model. We show that a training set restricted to data for a clean superconductor can correctly predict the topological phase diagram for a disordered superconductor. The accuracy of the ML prediction improves if the training set also includes disordered systems. In addition, we show that the imaginary part of the odd-frequency local pairing function shows the same spatial dependence as the LDOS, including signatures of edge states, and a simple linear regression model can predict the former from the latter.
@@
This paper is organized as follows. In Sec. II, we introduce the model and discuss trivial and topological edge states in the clean case. In Sec. III, we describe the phase diagram of the disordered system based on the calculation of the Bott index. In Sec. IV, we present our results for the ML prediction of the Bott index from the LDOS. Section V contains our results on the relation between odd-frequency pairing and the LDOS. Finally, we offer some concluding remarks in Sec. VI.
@@ II General model and clean case
We start by analyzing the edge modes in a clean system ($p=0$). For this purpose, we first consider a lattice with PBC in the $x$ direction and OBC in the $y$ direction. Using the translational invariance along $x$, we define
@@
where $k_{x}$ corresponds to the momentum in the $x$ direction. We then take the Fourier transform to Matsubara frequencies and the analytical continuation $i\omega\to\omega+i\eta$ to obtain the retarded Green’s function $G^{\rm ret}_{\nu\nu}(k_{x},y;\omega)$. In Fig. 1 we show the edge spectral function $A(k_{x},\omega)=-\frac{1}{\pi}\sum_{\nu=\uparrow,\downarrow}\text{Im}G^{\rm ret}_{\nu\nu}(k_{x},y=0;\omega)$ at the lower edge of the system. In Fig. 1(a), the chemical potential lies in the topological phase with $C=2$. As expected, we see two gapless dispersing edge states inside the bulk gap in this case. By contrast, gapless edge modes are absent in Fig. 1(b), corresponding to a value of chemical potential in the trivial phase.
@@ IV Predicting the Bott index from the LDOS
Figure 5: ML prediction trained with disordered systems. Here we trained an ML model using data for a $71\times 51$ grid in the $\mu\times p$ phase diagram of Fig. 3(a). The training set has $N_{d}=47$ realizations for each value of $\mu$ and $p$. With this model we predict $B$ for a validation set of $N_{d}=21$ new realizations for each $\mu$ and $p$. Panel (a) shows the predicted $B$ averaged over $N_{d}$. The solid (dashed) curves represent critical values of $p(\mu)$ for $B$ in Fig. 3. In panel (b) we fix $\mu=-t$ and count how often each value of $B$ appears. Here we used a neural network with two hidden layers with $600$ nodes each, rectified linear unit as activation function and 5 epochs. To allow any integer as output, we considered an exit layer with a single node, after which we round the result to the nearest integer. We use mean squared error as loss function. When choosing the training set, we excluded data points near the transition, in the region between solid and dashed or dot-dashed lines. The accuracy in the validation set is above 99.5% outside this transition region.
@@
Recall that Eq. (6) provides us with the critical chemical potentials where topological phase transitions take place in the clean superconductor. Therefore, we know $B$ for each point along the $p=0$ axis of the phase diagram in Fig. 3(a). Our first task is to use the LDOS of these clean superconductors with their corresponding $B$ [see Fig. 4(a) for an example] to train an ML model to predict $B$ for disordered superconductors with $p>0$. For this purpose, we use a neural network with no hidden layer to classify superconductors into three classes, corresponding to $B=-1,0,+2$. In Fig. 4(b), we show the disorder-averaged prediction for $B$ as a function of $\mu$ and $p$. For comparison, we also show lines extracted from the direct calculation of the Bott index in Fig. 3(a). The solid lines represent the values of $p$ above which every disorder realization has $B=0$, whereas below the dashed (dot-dashed) line in the left (right) dome every realization has $B=-1$ ($B=2$). Surprisingly, although this model was trained only at $p=0$ it can predict qualitatively well the average value of $B$ at finite disorder strength. In Fig. 4(c), we fix $\mu=-t$ and show how often each predicted value of the Bott index appears as a function of $p$. Since the value $B=+1$ does not occur in the training set for $p=0$, this classification model is not able to predict the $B=+1$ realizations seen in Fig. 3(b). Moreover, it incorrectly predicts $B=-1$ realizations near the transition between the topological phase with $B=2$ and the trivial phase. Thus, this method provides reasonable predictions for the average $B$ but not for the frequency of the possible values of $B$ near the topological transition.
@@
In order to obtain an ML method able to predict a Bott index not included in the training set $\{-1,0,+2\}$, we consider a neural network where the exit neural layer has a single node, which can predict any real number. We then round the result to the nearest integer. To improve the accuracy of the model, we consider two hidden layers of neurons. We include realizations with $p>0$ in the training set, but we avoid the region near the transition. The idea is to check whether this model finds the $B=+1$ realizations without the inclusion of any realization with this topological invariant in the training set. The result of this approach is shown in Fig. 5. Figure 5(a) shows the average of $B$ over disorder realizations. Clearly, this result is better than Fig. 4(b) since the predicted topological transitions now take place between the solid and dashed lines extracted from Fig. 3(a). The accuracy of this model outside the transition region is higher than $99.5\%$. Figure 5(b) shows the Bott indices counting for $\mu=-t$ as a function of $p$. Remarkably, this model predicts $B=+1$ realizations in the topological transition, in agreement with the direct calculation of the Bott index. We note, however, that the counting of predicted $B=+1$ here is only qualitatively similar to the counting shown in Fig. 3(b).
@@ V Odd-frequency pairing
We now turn to the anomalous Green’s function. Taking the Fourier transform of Eq. (5) and the analytical continuation $i\omega\to\omega+i\eta$ with $\eta\to 0^{+}$, we obtain the retarded anomalous Green’s function $F^{\rm ret}_{\nu_{1}\nu_{2}}(\mathbf{r}_{1},\mathbf{r}_{2};\omega)$. We focus on local correlations, $\mathbf{r}_{1}=\mathbf{r}_{2}=\mathbf{r}$, with $\nu_{1}=\downarrow$ and $\nu_{2}=\uparrow$. To simplify the notation, we define $F(\mathbf{r},\omega)=F^{\rm ret}_{\downarrow\uparrow}\big{(}\mathbf{r},\mathbf{r};\omega\big{)}$. Since this correlation is local, it is an even-parity $s$-wave correlation. The pairing function $F(\mathbf{r},\omega)$ may be decomposed into even-frequency singlet $F^{e}(\mathbf{r},\omega)$ and odd-frequency triplet $F^{o}(\mathbf{r},\omega)$ components in the form Perrin et al. (2020); Santos et al. (2020)
@@ V.1 Predicting the anomalous Green’s function from the LDOS
While the pairing function is not directly measured in experiments, it can be estimated using the LDOS in the case of a single magnetic impurity or dilute magnetic impurities in superconductors Perrin et al. (2020); Santos et al. (2020). Motivated by this fact, here we train a ML model of linear regression using $\rho(\mathbf{r},0)$ in order to obtain $F^{\prime\prime}(\mathbf{r},0)$. The training set consists of $N_{d}=47$ disorder realizations for each value of $\mu$ and $p$. The testing set has $N_{d}=21$ new realizations for each $\mu$ and $p$, for which we predict $F^{\prime\prime}(\mathbf{r},0)$ using only the corresponding LDOS. Comparing the predicted results with the actual functions, we find that the root-mean-square error of the set of all 21 realizations is $\delta=0.0083$. In Fig. 6 we show the LDOS, the ML prediction and the actual $F^{\prime\prime}(\mathbf{r},0)$ of a realization with $p=0.3$ and three different values of $\mu$ corresponding to $B=0,-1,2$. With the simple linear regression method used here, we conclude that the predicted functions have reasonable agreement with the actual odd-frequency pairings at zero energy. Neural networks with hidden neuron layers may be used in order to improve accuracy. However, since the exit layer would now have 900 nodes corresponding to each position of the $30\times 30$ lattice, higher computational resources would be necessary.
V.2 Predicting the Bott index from the the anomalous Green’s function
@@
In Fig. 7(a) we show the result obtained when we use a Random Forest model trained with the real part of the anomalous Green’s functions, $F^{\prime}(\mathbf{r},\omega=0)$. The result clearly fails because it incorrectly predicts that the topological phase with $B=-1$ extends all the way to $p=1$. We tried several other ML models, including neural networks, and all of them predicted topological phase diagrams in poor agreement with the result in Fig. 3(a). This is in contrast with the result discussed in Fig. 4(b), where ML trained with the LDOS of clean systems was enough to predict qualitatively well the phase diagram.

### v2

@@ I Introduction
In parallel to these developments, magnetic impurities in superconductors have been shown to generate odd-frequency pairing Kuzmanovski et al. (2020); Perrin et al. (2020); Santos et al. (2020); Suzuki et al. (2022). The study of odd-frequency superconductivity started when Berezinskii Berezinskii (1974) pointed out that the symmetry of the superconducting state allows for a two-fermion pairing function which is odd under time reversal but even under the product of spatial inversion and spin and orbital permutations. Since then, states exhibiting this type of unconventional pairing have been identified in many different systems; for reviews, see Refs. Tanaka et al. (2012); Linder and Balatsky (2019). For a single magnetic impurity or dilute magnetic impurities in superconductors, the local density of states (LDOS) measured by STS can be used to estimate the imaginary part of the odd-frequency local pairing function Perrin et al. (2020); Santos et al. (2020). Remarkably, odd-frequency pairing has also been linked to the presence of Majorana bound states Daino et al. (2012); Asano and Tanaka (2013); Kashuba et al. (2017); Tsintzis et al. (2019); Tanaka et al. (2024). In particular, the anomalous Green’s function near vortices in chiral $p$-wave superconductors shows the same spatial structure as the LDOS associated with subgap Majorana modes Daino et al. (2012).
@@
In this work, we study the relation between topological invariants, edge states, and odd-frequency pairing in a model for a 2D superconductor with randomly distributed magnetic impurities. Besides being naturally present in superconducting monolayers Ménard et al. (2019), the disorder can be introduced by substitution of magnetic adatoms and lattice reconstruction during the sample growth Goedecke et al. (2022) or by direct atomic manipulation Küster et al. (2022). Sufficiently strong disorder can drive transitions between topological and trivial phases Qin et al. (2016); Mascot et al. (2019). In Ref. Mascot et al. (2019), these transitions were analyzed using the Bott index. This topological invariant can be calculated in real space and is equivalent to the Chern number in the homogeneous case Hastings and Loring (2010); Loring and Hastings (2011); Hastings and Loring (2011); Toniolo (2022). The Bott index has been used to characterize topological phases in several disordered and amorphous systems Bandres et al. (2016); Agarwala and Shenoy (2017); Huang and Liu (2018); Focassio et al. (2021); Ghadimi et al. (2021); Dantas and Andrade (2022).
@@
We are interested in how the topological classification based on the Bott index correlates with the LDOS measured in STS experiments. To address this question, we use machine learning (ML) Carrasquilla (2020), a provably efficient method to classify phases of matter Huang et al. (2022); Canabarro et al. (2019); Che et al. (2024); Lewis et al. (2024); Rouzé et al. (2024). In fact, ML techniques have been successfully applied to identify topological phase transitions van Nieuwenburg et al. (2017); Deng et al. (2017); Zhang et al. (2018); Rem et al. (2019); Rodriguez-Nieva and Scheurer (2019); Scheurer and Slager (2020); Lustig et al. (2020); Claussen et al. (2020); Holanda and Griffith (2020); Che et al. (2020); Molignini et al. (2021); Margalit et al. (2022), to estimate model parameters from numerical or experimental data Wang et al. (2017); Gentile et al. (2021); Koch et al. (2023), and to analyze images generated by scanning probe techniques in quantum materials Burzawa et al. (2019); Iwasawa et al. (2022); Matthies et al. (2022); Basak et al. (2023); Sobral et al. (2023). Here we train ML models to predict the Bott index from the LDOS calculated numerically for our model. We show that a training set restricted to data for a clean superconductor can correctly predict the topological phase diagram for a disordered superconductor. The accuracy of the ML prediction improves if the training set also includes disordered systems. In addition, we show that the imaginary part of the odd-frequency local pairing function shows the same spatial dependence as the LDOS, including signatures of edge states, and a simple linear regression model can predict the former from the latter.
@@
This paper is organized as follows. In Sec. II, we introduce the model and discuss trivial and topological edge states in the clean case. In Sec. III, we describe the phase diagram of the disordered system based on the calculation of the Bott index. In Sec. IV, we present our results for the ML prediction of the Bott index from the LDOS. Section V contains our results on the relation between odd-frequency pairing and the LDOS. We offer some concluding remarks in Sec. VI. Finally, Appendix A contains some discussion about the weights used in the ML model and their relation to features in the LDOS.
@@ II General model and clean case
We start by analyzing the edge modes in a clean system with $p=0$. Note that here “clean” refers to the absence of disorder, not the absence of magnetic impurities. For this purpose, we first consider a lattice with PBC in the $x$ direction and OBC in the $y$ direction. Using the translational invariance along $x$, we define
@@
where $k_{x}$ corresponds to the momentum in the $x$ direction. We then take the Fourier transform to Matsubara frequencies and the analytical continuation $i\omega\to\omega+i\eta$ to obtain the retarded Green’s function $G^{\rm ret}_{\nu\nu}(k_{x},y;\omega)$. To analyze the spectrum of the finite-size system, we use small but finite values of $\eta$, corresponding to a phenomenological broadening as used, for instance, in Ref. Perrin et al. (2020). In Fig. 1 we show the edge spectral function $A(k_{x},\omega)=-\frac{1}{\pi}\sum_{\nu=\uparrow,\downarrow}\text{Im}G^{\rm ret}_{\nu\nu}(k_{x},y=0;\omega)$ at the lower edge of the system. In Fig. 1(a), the chemical potential lies in the topological phase with $C=2$. As expected, we see two gapless dispersing edge states inside the bulk gap in this case. By contrast, gapless edge modes are absent in Fig. 1(b), corresponding to a value of chemical potential in the trivial phase.
@@ III Bott index and disorder-induced topological transitions
Figure 5: ML prediction trained with disordered systems. Here we trained an ML model using data for a $71\times 51$ grid in the $\mu\times p$ phase diagram of Fig. 3(a). The training set has $N_{d}=47$ realizations for each value of $\mu$ and $p$. With this model, we predict $B$ for a validation set of $N_{d}=21$ new realizations for each $\mu$ and $p$. Panel (a) shows the predicted $B$ averaged over $N_{d}$ realizations. The solid and dashed lines are the same as in Fig. 4(b). In panel (b) we fix $\mu=-t$ and count how often each value of $B$ appears. In panel (c), we show multiple phase transitions by fixing $p=0.3$ and varying $\mu$. Here we used a neural network with two hidden layers with $600$ nodes each, rectified linear unit as activation function, and 5 epochs. To allow any integer as output, we considered an exit layer with a single node, after which we round the result to the nearest integer. We use mean squared error as loss function. When choosing the training set, we excluded data points near the transition, in the region between solid and dashed or dot-dashed lines. The accuracy in the validation set is above 99.5% outside this transition region.
@@ IV Predicting the Bott index from the LDOS
Recall that Eq. (6) provides us with the critical chemical potentials where topological phase transitions take place in the clean superconductor. Therefore, we know $B$ for each point along the $p=0$ axis of the phase diagram in Fig. 3(a). Our first task is to use the LDOS of these clean superconductors with their corresponding $B$ [see Fig. 4(a) for an example] to train an ML model to predict $B$ for disordered superconductors with $p>0$. For this purpose, we use a neural network with no hidden layer and three exit nodes. The exit layer is influenced by a softmax activation function, which causes the three exit neurons to provide probabilities that the data belongs to each of the three classes labeled by the Bott index. In the absence of hidden layers, the neural network attempts to learn a linear function that maps the LDOS to one of the three classes. By analyzing the weights that the neural network assigns to the LDOS for all lattice sites in different phases (see Appendix A), we notice that in the trivial phase the weights tend to be higher in the bulk, whereas the weights associated to positions along the edge become higher in the topological phases. We interpret these weights as a sign that the ML model learns to identify patterns in the edge states, which are then associated with the different values of the Bott index.
@@
In Fig. 4(b), we show the disorder-averaged prediction for $B$ as a function of $\mu$ and $p$, based on training the ML model with data along the $p=0$ line. For comparison, we also show lines extracted from the direct calculation of the Bott index in Fig. 3(a). The solid lines represent the values of $p$ above which every disorder realization has $B=0$, whereas below the dashed (dot-dashed) line in the left (right) dome every realization has $B=-1$ ($B=2$). Although this model was trained only at $p=0$, it can predict qualitatively well the average value of $B$ at finite disorder strength. This result is surprising given that disorder leads to complex spatial patterns that differ significantly from the clean case. In Fig. 4(c), we fix $\mu=-t$ and show how often each predicted value of the Bott index appears as a function of $p$. Since the value $B=+1$ does not occur in the training set for $p=0$, this classification model is not able to predict the $B=+1$ realizations seen in Fig. 3(b). Moreover, it incorrectly predicts $B=-1$ realizations near the transition between the topological phase with $B=2$ and the trivial phase. Thus, this method provides reasonable predictions for the average $B$ but not for the frequency of the possible values of $B$ near the topological transition.
@@
Next, to improve the accuracy of the model, we include realizations with $p>0$ in the training set, but we avoid the region near the transition. Moreover, we now consider a neural network with two hidden layers where the exit neural layer has a single node. The hidden layers enable the network to learn complex patterns and relationships between the input and output data. The single neuron without softmax activation function in the exit layer allows us to obtain any real number as output, which is then rounded to the nearest integer. This method is able to predict a Bott index not included in the training set $\{-1,0,+2\}$. The idea is to check whether this model finds the $B=+1$ realizations without the inclusion of any realization with this topological invariant in the training set. The result of this approach is shown in Fig. 5. Figure 5(a) shows the average of $B$ over disorder realizations. Clearly, this result is better than Fig. 4(b) since the predicted topological transitions now take place between the solid and dashed lines extracted from Fig. 3(a). The accuracy of this model outside the transition region is higher than $99.5\%$. Figure 5(b) shows the Bott indices counting for $\mu=-t$ as a function of $p$. In Fig. 5(c) we can see multiple phase transitions as we vary $\mu$ while fixing the value of $p$. Remarkably, this model predicts $B=+1$ realizations near the topological transition between $B=0$ and $B=+2$, in agreement with the direct calculation of the Bott index. We note, however, that the counting of predicted $B=+1$ here is only qualitatively similar to the counting shown in Fig. 3(b).
@@ V Odd-frequency pairing
We now turn to the anomalous Green’s function. Taking the Fourier transform of Eq. (5) and the analytical continuation $i\omega\to\omega+i\eta$, we obtain the retarded anomalous Green’s function $F^{\rm ret}_{\nu_{1}\nu_{2}}(\mathbf{r}_{1},\mathbf{r}_{2};\omega)$. We focus on local correlations, $\mathbf{r}_{1}=\mathbf{r}_{2}=\mathbf{r}$, with $\nu_{1}=\downarrow$ and $\nu_{2}=\uparrow$. To simplify the notation, we define $F(\mathbf{r},\omega)=F^{\rm ret}_{\downarrow\uparrow}\big{(}\mathbf{r},\mathbf{r};\omega\big{)}$. Since this correlation is local, it is an even-parity $s$-wave correlation. The pairing function $F(\mathbf{r},\omega)$ may be decomposed into even-frequency singlet $F^{e}(\mathbf{r},\omega)$ and odd-frequency triplet $F^{o}(\mathbf{r},\omega)$ components in the form Perrin et al. (2020); Santos et al. (2020)
@@ V.1 Predicting the anomalous Green’s function from the LDOS
While the pairing function is not directly measured in experiments, it can be estimated using the LDOS in the case of a single magnetic impurity or dilute magnetic impurities in superconductors Perrin et al. (2020); Santos et al. (2020). Motivated by this fact, here we train a ML model of linear regression using $\rho(\mathbf{r},0)$ to obtain $F^{\prime\prime}(\mathbf{r},0)$. The training set consists of $N_{d}=47$ disorder realizations for each value of $\mu$ and $p$. The testing set has $N_{d}=21$ new realizations for each $\mu$ and $p$, for which we predict $F^{\prime\prime}(\mathbf{r},0)$ using only the corresponding LDOS. Comparing the predicted results with the actual functions, we find that the root-mean-square error of the set of all 21 realizations is $\delta=0.0083$. In Fig. 6 we show the LDOS, the ML prediction and the actual $F^{\prime\prime}(\mathbf{r},0)$ of a realization with $p=0.3$ and three different values of $\mu$ corresponding to $B=0,-1,2$. With the simple linear regression method used here, we conclude that the predicted functions have reasonable agreement with the actual odd-frequency pairings at zero energy. Neural networks with hidden neuron layers may be used in order to improve accuracy. However, since the exit layer would now have 900 nodes corresponding to each position of the $30\times 30$ lattice, higher computational resources would be necessary.
V.2 Predicting the Bott index from the anomalous Green’s function
@@
In Fig. 7(a) we show the result obtained when we use a Random Forest model trained with the real part of the anomalous Green’s functions, $F^{\prime}(\mathbf{r},\omega=0)$. The result clearly fails because it incorrectly predicts that the topological phase with $B=-1$ extends all the way to $p=1$. We tried several other ML models, including neural networks, and all of them predicted topological phase diagrams in poor agreement with the result in Fig. 3(a). This is in contrast with the result discussed in Fig. 4(b), where ML trained with the LDOS of clean systems was enough to predict qualitatively well the phase diagram. To interpret this result, we note that the real and imaginary parts of (normal or anomalous) Green’s functions are related by Kramers-Kronig relations. However, these relations involve an integral over the full energy spectrum. For this reason, the real part of the anomalous Greens function $F^{\prime}(\mathbf{r},\omega=0)$ is influenced by the imaginary part $F^{\prime\prime}(\mathbf{r},\omega)$ for every energy $\omega$. As a consequence, while $F^{\prime\prime}(\mathbf{r},\omega=0)$ shows clear signatures of topological edge states, we do not observe these signatures in $F^{\prime}(\mathbf{r},\omega=0)$. This explains why the prediction with the real part in Fig.  7(a) is so poor. It also indicates that the main feature that the neural network uses to classify the images is the enhancement of the LDOS near the edges in the topological phases.